• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name LeRose

    Log entry time 16:04:06 on April 11, 2007

    Entry number 198638

    Followups:

    keyword=Shift Summary

    John LeRose SL
    Ron Gilman TO

    Kinematics Q^2=55fm^-2
    Left: 0.5964GeV/c 140.301°
    Right: 0.72620GeV/c 14.945°

    12:15 DAQ problems ROC3 won't reset. Bob fixed it
    (lost memory of what it needed to reboot)

    13:07 MCC takes beam for injector work
    14:20 Left Q1 tripped. Reset and it came right back.
    16:00 still waiting for beam

    runs:
    2320 07:51-08:13 3He 1.4M 70μA
    2321 08:14-08:43 3He 1.4M 70μA
    2322 08:44-09:07 3He 1.4M 70μA
    2323 09:07-09:31 3He 1.4M 70μA
    2324 09:32-10:00 3He 1.4M 70μA
    2325 10:01-10:25 3He 1.4M 70μA
    2326 10:25-10:52 3He 1.4M 70μA
    2327 10:53-11:16 3He 1.4M 70μA
    2328 11:16-11:42 3He 1.4M 70μA
    2329 11:42- DAQ problem
    2330 12:48-13:08 3He 0.708M 70μA

    Note: See halog 198634. Radiation monitor strip chart is avaliable
    to monitor beam quality.